A new drug cures 9 of 200 patients suffering from a type of cancer, for which the historical cure rate is 2%. Perform a test to check on the significance of this result, at both 5% and 1% levels of significance. Based on your conclusion, comments on the efficacy of the treatment using this new drug.


1
Expert's answer
2021-05-03T05:04:05-0400

One Proportion z-test

"p_0 = 2 \% = 0.02 \\

H_0 : p = 0.02 \\

H_1 : p > 0.02 \\

n = 200"

The number of successes:

x = 9

The sample proportion:

"\hat{p} = \frac{x}{n}= \frac{9}{200}= 0.045"

The test statistics:

"z = \frac{\hat{p} -p_0}{ \sqrt{ \frac{p_0(1-p_0)}{n} }} \\

= \frac{0.045 -0.02}{ \sqrt{ \frac{0.02 \times (1-0.02)}{200} }} \\

= 2.5254 \\

P(z> 2.5254) = 1 -P(Z<2.5254) = 0.0058"

At the 5 % significance level (α=0.05), we reject the H0, since p-value < α. We have sufficient evidence to conclude that the new drug has a better efficacy than the historical cure rate in curing the type of cancer at 5 % significance level.

At the 1 % significance level (α=0.01), we reject the H0, since p-value < α. We have sufficient evidence to conclude that the new drug has a better efficacy than the historical cure rate in curing the type of cancer at 1 % significance level.